OpenStudy (anonymous):

f(x)=x+7/2X-1 domain?

OpenStudy (john_es):

I suppose, \[f(x)=\frac{x+7}{2x-1}\]In order to obtain the domain you must remember that you cannot divide by zero, so, \[2x-1=0\Rightarrow x=1/2\]Then \[D=\mathbb{R}-\{1/2\}\]